A box contains 2 black balls and 3 gold balls. 2 balls are randomly drawn in succession from the box.
A. If there is no replacement, what is the probability that both balls are black?
B. If there is replacement before the secon draw, what is the probability that both balls are black?
Thanks!

Answer by rfer(2688)

A) 2/5*1/4=2/20=1/10
B) 2/5*2/5=4/25